If P(A | B) = 0.2, P(A) = 0.3 and P(B) = 0.4, what is P(A or B) ?
Answer by ikleyn(52908)

.

By the definition of the conditional probability,


    P(A | B) = P(A ∩ B)/P(B).


Therefore, by substituting given data  P(A | B) = 0.2  and P(B) = -/4


    0.2 = P(A ∩ B)/0.4,


which implies  P(A ∩ B) = 0.2*0.4 = 0.08.


Next,  P(A U B) = P(A) + P(B) - P(A ∩ B) = 0.3 + 0.4 - 0.08 = 0.62.   ANSWER

Completed and solved.

------------------

On conditional probability, see the lesson
    - Conditional probability problems
in this site.